CocoCammron:

5(- 3x - 2) - (x - 3) = -4(4x + 5) + 13

supie:

Do I just solve it, @CocoCammron?

CocoCammron:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Do I just solve it, @CocoCammron? \(\color{#0cbb34}{\text{End of Quote}}\) yeah

umm:

Is there answer choices?

CocoCammron:

No

umm:

Is there a way you can show what it's asking, please?

supie:

I think we're just solving \(5(- 3x - 2) - (x - 3) = -4(4x + 5) + 13\)

sddqdqd:

what did you get and i can tell you if you were correct

CocoCammron:

It is a homework assiment THere is no answer choices

sddqdqd:

Infinitely Many Solutions

umm:

\(\color{#ffff4d}{\text{Originally Posted by}}\) @sddqdqd Infinitely Many Solutions \(\color{#ffff4d}{\text{End of Quote}}\) Please provide how you got your answer.

supie:

\(5(−3x−2)−(x−3)=−4(4x+5)+13\) Simplify, both sides and distribute the`-` sign, \(5(−3x−2)+−1(x−3)=−4(4x+5)+13\) \(5(−3x−2)+−1x+(−1)(−3)=−4(4x+5)+13\) \(5(−3x−2)+−x+3=−4(4x+5)+13\) Distribute: \((5)(−3x)+(5)(−2)+−x+3=(−4)(4x)+(−4)(5)+13\) \(−15x+−10+−x+3=−16x+−20+13\) Combine the like terms: \((−15x+−x)+(−10+3)=(−16x)+(−20+13)\) \(−16x+−7=−16x+−7\) \(−16x−7=−16x−7\) \(+16x\) to both sides \(−16x−7+16x=−16x−7+16x=\) \(−7+7=−7+7\) \(0=0\), true

sddqdqd:

^ i got mine the same way

umm:

However, that would only mean all real numbers are solutions.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @umm However, that would only mean all real numbers are solutions. \(\color{#0cbb34}{\text{End of Quote}}\) I know. Thanks.

umm:

Yes, I was explaining to @sddqdqd since their answer was infinitely many solutions.

sddqdqd:

since both sides equal, there are infinitely many solutions.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @umm Yes, I was explaining to @sddqdqd since their answer was infinitely many solutions. \(\color{#0cbb34}{\text{End of Quote}}\) Oh right.